OpenStudy (anonymous):

so find for me the derivertive of thise one 3x^2+siny^2+y=e^4x

OpenStudy (anonymous):

derivative with respect to ??? (x or y)

OpenStudy (anonymous):

with respect to x @PATTOOO

OpenStudy (anonymous):

6x=4*e^4x other y terms will be 0

OpenStudy (anonymous):

thise is an implicit diffriatiation hence findin dy/dx can be found differently and y can not be zero

OpenStudy (anonymous):

ohh sry

OpenStudy (anonymous):

\[\sin(y^{2}) = e^{4x}-3x^{2} \] \[2 y cos(y^{2}) \frac{dy}{dt}=4e^{4x}\frac{dx}{dt}-6x\frac{dx}{dt}\] \[\frac {dy} {dt}(2ycos(y^2)) = \frac{dx}{dt}(4e^{4x}-6x)\] \[y'=\frac{(2e^{4x}-3x)sec(y^2)}{y}\]